OpenStudy (anonymous):

The volumes of two similar are 1,331 m^3 and 216m^3. The surface area of the larger solid is 484m^2. What is the surface area of the smaller solid? 864m^2 288m^2 144m^2 68m^2

OpenStudy (anonymous):

volume ratio =1331/216=6.16

OpenStudy (anonymous):

@AJO1 what do I do next to get the surface area??

OpenStudy (anonymous):

@AJ01

OpenStudy (anonymous):

Idk cause that's not one of my answers that were given @AJ01

OpenStudy (anonymous):

it should ......

OpenStudy (anonymous):

the scalar factor should be the same

OpenStudy (anonymous):

Your right but idk what to do @AJ01

OpenStudy (anonymous):

i will i am not sure. it con't be A

OpenStudy (anonymous):

Yeah I know

OpenStudy (anonymous):

Could it be B ?? @AJ01

OpenStudy (anonymous):

1331/484=216/x x=216*484/1331=78.5 -------- x=484/6.16=78.5

OpenStudy (anonymous):

B is half of it i would go with D